What dienes containing isolated double bonds are c
Answer #1

CH 3 C-CH-CH=CH OH CH3_ CH OH H H Cxt 2 CH CH uolatal diane stabla) A Heuce 4 capable CH CH lH CH2= CH-CH=C-CHI-CH3 uolated d